OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y = 2x^2 and y = 3 and x = 0 about the y-axis.

OpenStudy (anonymous):

i think this is just \[\pi\int_0^3\frac{y}{2}dy\] malevolence?

OpenStudy (anonymous):

Your region is only the left side. http://www.wolframalpha.com/input/?i=plot+y%3D2x^2%2Cy%3D3 Set 2x^2=3 x=+/-sqrt(3/2) Use 0 to the positive one. Using shell method: The height is 3-2x^2; the radius is x You have: \[2 \pi \int\limits_0^{\frac{\sqrt3}{2}} x(3-2x^2)dx\] Which gives \[\frac{9 \pi}{4}\]

OpenStudy (anonymous):

Should read "region is on right side"

OpenStudy (anonymous):

sole for x and get \[x=\sqrt{\frac{y}{2}}\] then use \[\int \pi r^2\] yes?

OpenStudy (anonymous):

hold the phone. you are rotating around y - axis yes?

OpenStudy (anonymous):

Yes, and you are given y=2x^2 y=3 so you should use shell method so you can slice parallel to the axis of rotation.

OpenStudy (anonymous):

satellite, we get the same answer, you just used a different method :P

OpenStudy (anonymous):

yes or you can use \[\pi r^2\] get the same thing. i think integrating wrt y is easiest here maybe

OpenStudy (anonymous):

you are right. get the same thing

OpenStudy (anonymous):

You're probably right, I just saw a x^2 and was like "solving for x gives a pm" so I went with shell

OpenStudy (anonymous):

got it. here is a picture of the region to be rotated. that is why i used that method http://www.wolframalpha.com/input/?i=y+%3D+2x^2+%2C+y+%3D+3+%2C+